# Random Pareto front surfaces

### v1

@@ Abstract
The Pareto front of a set of vectors is the subset which is comprised solely of all of the best trade-off points. By interpolating this subset, we obtain the optimal trade-off surface. In this work, we prove a very useful result which states that all Pareto front surfaces can be explicitly parametrised using polar coordinates. In particular, our polar parametrisation result tells us that we can fully characterise any Pareto front surface using the length function, which is a scalar-valued function that returns the projected length along any positive radial direction. Consequently, by exploiting this representation, we show how it is possible to generalise many useful concepts from linear algebra, probability and statistics, and decision theory to function over the space of Pareto front surfaces. Notably, we focus our attention on the stochastic setting where the Pareto front surface itself is a stochastic process. Among other things, we showcase how it is possible to define and estimate many statistical quantities of interest such as the expectation, covariance and quantile of any Pareto front surface distribution. As a motivating example, we investigate how these statistics can be used within a design of experiments setting, where the goal is to both infer and use the Pareto front surface distribution in order to make effective decisions. Besides this, we also illustrate how these Pareto front ideas can be used within the context of extreme value theory. Finally, as a numerical example, we applied some of our new methodology on a real-world air pollution data set.
@@ 1 Introduction
The Pareto front is often regarded as the natural generalisation of the maximum (or minimum) for vector-valued sets. This generalisation is based on the Pareto partial ordering relation, which gives us a way to compare between any two vectors when possible. Geometrically, the Pareto front can be used to define a surface in the vector space which describes the best possible trade-offs one can obtain between the different components. The primary focus of this work will be to study this space of Pareto front surfaces in more detail. To accomplish this, we establish our main result in Theorem 3.1, which gives us an explicit way to parametrise any Pareto front surface using polar coordinates. The rest of the work is then focussed on using this polar parametrisation result in order to generalise existing concepts from linear algebra, probability and statistics, and decision theory to function over this exotic space of Pareto front surfaces. In particular, we focus our attention on the stochastic setting where the Pareto front surface itself is random. Under this regime, our polar parametrisation result tells us that we can treat any random Pareto front surface as an infinite-dimensional stochastic process (10), which is indexed by the set of positive unit vectors. This stochastic process can then be used along with standard ideas from functional data analysis in order to define and estimate any Pareto front surface statistic of interest in a principled and efficient manner—more details in Section 4. We then proceed to show how these ideas can be used within a standard decision making workflow, where one is interested in actively identifying decisions which lead to the best trade-off for the decision maker—see Section 5.1 and Section 5.2. As another independent application, we also highlight how our work on random Pareto front surfaces can be applied within the domain of multivariate extreme value theory—more details in Section 5.3. Lastly, in Section 5.4, we present a numerical example which demonstrates how our ideas can be used in order to analyse Pareto fronts in time series data. Precisely, we focus our attention on the change in the maximum amount of daily air pollution in North Kensington, London. We find that the daily Pareto front of pollutant levels has indeed been reduced after the introduction of many recent pollution reduction initiatives.
1.1 Structure of the paper
The remainder of the paper is organised as follows: In Section 2, we introduce the key definitions and nomenclature that will be used throughout the paper. In Section 3, we present our main result, which gives an explicit parametrisation of any Pareto front surface based on polar coordinates. We then use this polar parametrisation in order to define some useful operations and concepts on the space of Pareto front surfaces. Notably we define the concept of an order-preserving transformation (Section 3.1), a length-based utility function (Section 3.2) and a length-based loss function (Section 3.3). In Section 4, we study the setting where the Pareto front surface is random and define many useful statistics of interest such as the expectation (Section 4.1), covariance (Section 4.2) and quantiles (Section 4.3), among others. We also connected our discussion with existing work on Pareto front surface distributions, which are based solely on ideas from random set theory (Section 4.6). In Section 5, we demonstrate the effectiveness of these new ideas on some practical applications. For instance, in Section 5.1, we present a novel visualisation strategy based our polar parametrisation result. In particular, we show how it is possible to construct a picture of the whole Pareto front surface by using a family of low-dimensional slices. Consequently, in Section 5.2, we illustrate how all of these uncertainty quantification and visualisation ideas can be used within a standard Bayesian experimental design setting. Specifically, we demonstrate how these tools can be used both during an optimisation routine (Section 5.2.1) and also in the post-decision setting (Section 5.2.2). The former stage focusses on the problem of identifying the best experiments to run, whilst the latter stage is focussed on identifying the inputs which are most likely going lead to the desired outputs. Subsequently, in Section 5.3, we demonstrate how it is possible to adapt some prominent results from extreme value theory to work in the multivariate setting, where the maximum is defined using the Pareto partial ordering. This is in contrast to most existing work in multivariate extreme value theory, which has largely focussed on the setting where the vector-valued maximum is defined in a component-wise fashion. Afterwards, in Section 5.4, we present a case study on how some of these Pareto front ideas can be used in order to quantify changes in the daily maximum air pollutant levels in a part of west London. In Section 6, we conclude this work and include a discussion on future research directions. Finally, in Appendix A, we include the proofs of all of the results that are stated within the paper. The code that is needed to reproduce the figures and numerical experiments are available in our Github repository: https://github.com/benmltu/scalarize.
@@ 3 Polar parametrisation
Notably we present an explicit representation for this isomorphism in (5), which we refer to as the polar parametrisation of a Pareto front surface. The name of this representation is motivated by the fact that we rely on the hyperspherical polar coordinates transformation in order to derive this result. The intuitive idea behind this result is concisely described in Figure 2. Informally speaking, we can identify each Pareto optimal point by drawing a line from the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$ along a positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. If we know the length of these lines $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}\in\mathbb{R}$ when it intersects the Pareto front surface, then we can reconstruct the Pareto front surface by using a linear mapping: $\boldsymbol{\eta}+\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}\boldsymbol{\lambda}\in\mathbb{R}^{M}$ for $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. In the following paragraphs, we will formalise this idea more concretely and give an explicit construction for this length function.
@@
To prove Theorem 3.1, we rely on the following coordinate system transformation $\mathcal{T}_{\boldsymbol{\eta}}:\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]\rightarrow\mathcal{S}^{M-1}_{+}\times\mathbb{R}_{>0}$,
@@
which is a non-negative [^3] The length scalarisation function is positive on the truncated space $\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$ and zero everywhere else. function that is defined for all vectors $\mathbf{y}\in\mathbb{R}^{M}$. Conceptually, when the objective vector lies in the truncated space $\mathbf{y}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$, then the length scalarisation function computes the length of the line $L_{\boldsymbol{\eta},\boldsymbol{\lambda}}:=\{\boldsymbol{\eta}+t\boldsymbol{\lambda}:t\in\mathbb{R}\}$ lying in the open hyper-rectangle $(\boldsymbol{\eta},\mathbf{y}):=\{\mathbf{z}\in\mathbb{R}^{M}:z^{(m)}\in(\eta^{(m)},y^{(m)})\text{ for }m=1,\dots,M\}$—see the left of Figure 4 for an illustration of this intuition in two dimensions. Consequently, the optimal direction function $\boldsymbol{\lambda}_{\boldsymbol{\eta}}^{*}:\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]\rightarrow\mathcal{S}^{M-1}_{+}$ is the function that returns the largest projected length
@@
Figure 2: An illustration of the polar parametrisation result in $M=2$ dimensions.
@@
$A\in\mathbb{L}_{\boldsymbol{\eta}}\iff A\subset\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]\cup\{\boldsymbol{\eta}\}\enskip\text{ and }\enskip|A\cap L_{\boldsymbol{\eta},\boldsymbol{\lambda}}|=1\text{ for all }\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}.$
Moreover, for this set of polar surfaces $\mathbb{L}_{\boldsymbol{\eta}}$, we define the projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}:\mathbb{L}_{\boldsymbol{\eta}}\rightarrow\mathbb{R}_{\geq 0}$ as the function which returns the $L^{2}$-distance of the intersected point along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$:
@@
Figure 3: An illustration of the polar parametrisation associated with a set of two points.
Note that there is a duality between a polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$ and its projected lengths:
$A\in\mathbb{L}_{\boldsymbol{\eta}}\longleftrightarrow\{\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[A]\geq 0:\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}\}.$
By design, we can easily reconstruct the whole polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$, from its collection of projected lengths, by using the equation
$A=\{\boldsymbol{\eta}+\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[A]\boldsymbol{\lambda}\in\mathbb{R}^{M}:\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}\}.$ (4)
@@
Theorem 3.1 (Polar parametrisation)
For any bounded set of vectors $A\subset\mathbb{R}^{M}$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if the corresponding Pareto front surface is not empty $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\neq\emptyset$, then it admits the following polar parametrisation:
@@
The explicit form of the polar parametrisation (5) can be derived in an constructive manner. Firstly, one can show that the Pareto front surface of a single point $\mathbf{a}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$ can be written as the polar surface
@@
Remarkably, this overall procedure turns out to be equivalent to just retaining the points which achieve the maximal projected lengths along each positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. An illustration of this conceptual idea is presented in Figure 3. Note however that the proof of the result in Section A.1 does not actually follow this intuitive construction. Instead, our proof is derived in a more technical way by taking advantage of a well-known optimisation result regarding the Chebyshev scalarisation function (Miettinen, 1998, Part 2, Theorem 3.4.5). The primary reason why we take this more general approach is to ensure that our polar parametrisation result holds for sets lying in the whole objective space $\mathbb{R}^{M}$ as opposed to just the truncated objective space $\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$.
@@
The coordinate system transformation that we performed in (1) is appealing because it admits a representation for the Pareto front surface, which is simple, explicit and has a linear dependency on the projected length values. Nevertheless, we can easily obtain other nonlinear representations of a Pareto front surface by taking advantage of transformation functions. For example, let $\tau:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ denote an invertible and strictly monotonically increasing transformation function such that $\mathbf{y}\succeq\mathbf{y}^{\prime}\iff\tau(\mathbf{y})\succeq\tau(\mathbf{y}^{\prime})$ for any $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{M}$. Then, by a simple monotonicity argument, we can show that the following parametrisation of the Pareto front surface is also valid:
@@
Intuitively, this reformulation works by performing the polar parametrisation in the transformed space before inverting the result. Note that strict monotonicity is required here because we want the Pareto front surface to be invariant under these transformations: $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]=\tau^{-1}(\mathcal{Y}_{\tau(\boldsymbol{\eta})}^{\textnormal{int}}[\tau(A)])$.
@@
For a finite set $A\subset\mathbb{R}^{M}$, the worst-case cost involved with computing the strictly Pareto optimal front $\mathcal{Y}^{\textnormal{strict}}[A]\subset\mathbb{R}^{M}$ is $\mathcal{O}(|A|^{2}M)$. In contrast, the worst-case cost involved with computing a finite approximation of the polar parametrised Pareto front surface $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\subset\mathbb{R}^{M}$ is $\mathcal{O}(|\Lambda||A|M)$, where $\Lambda=\{\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2},\dots\}$ denotes a finite set of positive unit vectors.

### v2

@@ Abstract
The goal of multi-objective optimisation is to identify the Pareto front surface which is the set obtained by connecting the best trade-off points. Typically this surface is computed by evaluating the objectives at different points and then interpolating between the subset of the best evaluated trade-off points. In this work, we propose to parameterise the Pareto front surface using polar coordinates. More precisely, we show that any Pareto front surface can be equivalently represented using a scalar-valued length function which returns the projected length along any positive radial direction. We then use this representation in order to rigorously develop the theory and applications of stochastic Pareto front surfaces. In particular, we derive many Pareto front surface statistics of interest such as the expectation, covariance and quantiles. We then discuss how these can be used in practice within a design of experiments setting, where the goal is to both infer and use the Pareto front surface distribution in order to make effective decisions. Our framework allows for clear uncertainty quantification and we also develop advanced visualisation techniques for this purpose. Finally we discuss the applicability of our ideas within multivariate extreme value theory and illustrate our methodology in a variety of numerical examples, including a case study with a real-world air pollution data set.
@@ 1 Introduction
The Pareto front is often regarded as the natural generalisation of the maximum (or minimum) for vector-valued sets. This generalisation is based on the Pareto partial ordering relation, which gives us a way to compare between any two vectors when possible. Geometrically, the Pareto front can be used to define a surface in the vector space which describes the best possible trade-offs one can obtain between the different components. Capturing this surface is crucial for many real-world applications and machine learning tasks where one is interested in jointly maximising multiple reward criteria. Notably, this problem becomes much more challenging when there is uncertainty in the rewards, which is very common in practice. The purpose of this work is to address this practically important problem directly by providing the tools and methodology required to study and effectively use Pareto front surfaces in the face of uncertainty. To the best of our knowledge, a study of this kind is currently missing in the literature and would be very useful for real-world practitioners who deal with these problems in practice.
At a high level, our novel framework is based on specific polar representation of the Pareto front surface. More precisely, we show that any Pareto front surface can be equivalently characterised by its projected length function, which is a scalar-valued function that computes the projected lengths of the polar surface along any positive radial direction. The clear benefit of this representation is that it is explicit and gives us an interpretable and efficient way to work over the space of Pareto front surfaces. To demonstrate the value of our framework, we focus our attention on the stochastic setting where the Pareto front surface itself is unknown and random. Under this regime, our polar parameterisation result allows us to treat any random Pareto front surface as a tractable infinite-dimensional random variable. Various statistics such as its expectation, covariance and quantiles are then well defined and straightforward to compute numerically. Importantly, these statistics are especially useful in the context of quantifying uncertainty and making better decisions in the stochastic setting. Overall, we believe that our methodology is very practical as it can be easily integrated into most standard decision making workflows and there are many potential directions to extend our work even further.
1.1 Structure and contributions
The remainder of the paper is organised as follows: In Section 2, we introduce the key definitions and nomenclature that will be used throughout the paper. In Section 3, we present our framework and show that any Pareto front surface can be expressed using polar coordinates. We then use this polar parameterisation in order to define some useful operations and concepts on the space of Pareto front surfaces. Namely we define the concept of an order-preserving transformation (Section 3.1), a length-based utility function (Section 3.2) and a length-based loss function (Section 3.3). We note that whilst polar representations have been used before as a technical tool for manipulating calculations related to hypervolume indicator (Shang et al., 2018; Deng and Zhang, 2019; Zhang and Golovin, 2020), this is the first time to the best of our knowledge that they have been studied in detail and proposed as a methodological tool.
In Section 4, we study the setting where the Pareto front surface is random and define many useful statistics of interest such as the expectation (Section 4.1), covariance (Section 4.2) and quantiles (Section 4.3), among others. We also discuss links with existing works on Pareto front surface distributions based on ideas from random set theory (Section 4.6). We then proceed to Section 5, where we show how these ideas can be used in practice:
  • In Section 5.1, we present a novel visualisation strategy based our polar parameterisation and we show how it is possible to construct a picture of the whole Pareto front surface by using a family of low-dimensional slices. This can assist in situations where one is interested in actively identifying decisions or inputs which lead to the best trade-offs for the decision maker.
  • In Section 5.2, we illustrate how all of these uncertainty quantification and visualisation ideas can be used within a standard Bayesian experimental design setting. Specifically, we demonstrate how these tools can be used both during an optimisation routine (Section 5.2.1) and also in the post-decision setting (Section 5.2.2). The former stage focusses on the problem of identifying the best experiments to run, whilst the latter stage is focussed on identifying the inputs which are most likely going lead to the desired outputs.
  • In Section 5.3, we demonstrate how it is possible to adapt some prominent results from extreme value theory to work in the multivariate setting, where the maximum is defined using the Pareto partial ordering. This is in contrast to most existing work in multivariate extreme value theory, which has largely focussed on the setting where the vector-valued maximum is defined in a component-wise fashion.
  • In Section 5.4, we present a case study on how some of these Pareto front ideas can be used in order to quantify changes in the daily maximum air pollutant levels in a part of west London.
In Section 6, we conclude this work and include a discussion on future research directions. Finally, in Appendix A, we include the proofs of all of the results that are stated within the paper. The code that is needed to reproduce the figures and numerical experiments are available in our Github repository: https://github.com/benmltu/scalarize.
@@ 3 Polar parameterisation
Notably we present an explicit representation for this isomorphism in (5), which we refer to as the polar parameterisation of a Pareto front surface. The name of this representation is motivated by the fact that we rely on the hyperspherical polar coordinates transformation in order to derive this result. The intuitive idea behind this result is concisely described in Figure 2. Informally speaking, we can identify each Pareto optimal point by drawing a line from the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$ along a positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. If we know the length of these lines $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}\in\mathbb{R}$ when it intersects the Pareto front surface, then we can reconstruct the Pareto front surface by using a linear mapping: $\boldsymbol{\eta}+\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}\boldsymbol{\lambda}\in\mathbb{R}^{M}$ for $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. In the following paragraphs, we will formalise this idea more concretely and give an explicit construction for this length function.
@@
To derive our polar parameterisation result, we rely on the following coordinate system transformation $\mathcal{T}_{\boldsymbol{\eta}}:\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]\rightarrow\mathcal{S}^{M-1}_{+}\times\mathbb{R}_{>0}$,
@@
which is a non-negative [^3] The length scalarisation function is positive on the truncated space $\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$ and zero everywhere else. function that is defined for all vectors $\mathbf{y}\in\mathbb{R}^{M}$. Conceptually, when the objective vector lies in the truncated space $\mathbf{y}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$, then the length scalarisation function computes the length of the line $L_{\boldsymbol{\eta},\boldsymbol{\lambda}}:=\{\boldsymbol{\eta}+t\boldsymbol{\lambda}:t\in\mathbb{R}\}$ lying in the open hyper-rectangle $(\boldsymbol{\eta},\mathbf{y})\subset\mathbb{R}^{M}$—see the left of Figure 3 for an illustration of this intuition in two dimensions. Consequently, the optimal direction function $\boldsymbol{\lambda}_{\boldsymbol{\eta}}^{*}:\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]\rightarrow\mathcal{S}^{M-1}_{+}$ is the function that returns the largest projected length
@@
Figure 2: An illustration of the polar parameterisation result in $M=2$ dimensions.
@@
$A\subset\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]\cup\{\boldsymbol{\eta}\}\enskip\text{ and }\enskip|A\cap L_{\boldsymbol{\eta},\boldsymbol{\lambda}}|=1\text{ for all }\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+},$
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. Moreover, for this set of polar surfaces $\mathbb{L}_{\boldsymbol{\eta}}$, we define the projected length function $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}:\mathbb{L}_{\boldsymbol{\eta}}\rightarrow\mathbb{R}_{\geq 0}$ as the function which returns the $L^{2}$-distance of the intersected point along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$:
@@
Note that the projected lengths provide a dual representation of a polar surface, that is
$A\in\mathbb{L}_{\boldsymbol{\eta}}\iff A=\{\boldsymbol{\eta}+\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[A]\boldsymbol{\lambda}\in\mathbb{R}^{M}:\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}\}.$ (4)
@@
Theorem 3.1 (Polar parameterisation)
For any bounded set of vectors $A\subset\mathbb{R}^{M}$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, if the corresponding Pareto front surface is not empty $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\neq\emptyset$, then it admits the following polar parameterisation:
@@
The explicit form of the polar parameterisation (5) can be derived in an constructive manner. Firstly, one can show that the Pareto front surface of a single point $\mathbf{a}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$ can be written as the polar surface
@@
Remarkably, this overall procedure turns out to be equivalent to just retaining the points which achieve the maximal projected lengths along each positive direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. An illustration of this conceptual idea is presented in Figure 4. Note however that the proof of the result in Section A.1 does not actually follow this intuitive construction. Instead, our proof is derived in a more technical way by taking advantage of a well-known optimisation result regarding the Chebyshev scalarisation function (Miettinen, 1998, Part 2, Theorem 3.4.5). The primary reason why we take this more general approach is to ensure that our polar parameterisation result holds for sets lying in the whole objective space $\mathbb{R}^{M}$ as opposed to just the truncated objective space $\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$.
@@
The coordinate system transformation that we performed in (1) is appealing because it admits a representation for the Pareto front surface, which is simple, explicit and has a linear dependency on the projected length values. Nevertheless, we can easily obtain other nonlinear representations of a Pareto front surface by taking advantage of transformation functions. For example, let $\tau:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ denote an invertible and strictly monotonically increasing transformation function such that $\mathbf{y}\succeq\mathbf{y}^{\prime}\iff\tau(\mathbf{y})\succeq\tau(\mathbf{y}^{\prime})$ for any $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{M}$. Then, by a simple monotonicity argument, we can show that the following parameterisation of the Pareto front surface is also valid:
@@
Intuitively, this reformulation works by performing the polar parameterisation in the transformed space before inverting the result. Note that strict monotonicity is required here because we want the Pareto front surface to be invariant under these transformations: $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]=\tau^{-1}(\mathcal{Y}_{\tau(\boldsymbol{\eta})}^{\textnormal{int}}[\tau(A)])$.
Figure 4: An illustration of the polar parameterisation associated with a set of two points.
@@
For a finite set $A\subset\mathbb{R}^{M}$, the worst-case cost involved with computing the strictly Pareto optimal front $\mathcal{Y}^{\textnormal{strict}}[A]\subset\mathbb{R}^{M}$ is $\mathcal{O}(|A|^{2}M)$. In contrast, the worst-case cost involved with computing a finite approximation of the polar parameterised Pareto front surface $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]\subset\mathbb{R}^{M}$ is $\mathcal{O}(|\Lambda||A|M)$, where $\Lambda=\{\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2},\dots\}$ denotes a finite set of positive unit vectors.